OpenStudy (broskishelleh):

Jesse kicks a soccer ball off the ground and in the air, with an initial velocity of 29 feet per second. Using the formula H(t) = -16t2 + vt + s, what is the maximum height the soccer ball reaches? 13.1 feet 13.5 feet 13.8 feet 14.2 feet

OpenStudy (anonymous):

+s? can you recheck the given forumla?

OpenStudy (broskishelleh):

Yep, this it @study100

OpenStudy (broskishelleh):

@study100 Please help!

OpenStudy (anonymous):

Alright, you can solve this by plug t= 0 and h(t) = 0 to the equation height = 0 at time = 0 and they say that initial velocity is 29. v= 29 0= 0 + 0 +s s= 0 now you can eliminate s out from the equation we just have H(t) = -16t^2+ vt

OpenStudy (broskishelleh):

Alright

OpenStudy (anonymous):

What math level is this? then you plug v=29 into the equation H(t) = -16t^2 +29t Now look at NUMBER 1 on this: http://www.wikihow.com/Find-the-Maximum-or-Minimum-Value-of-a-Quadratic-Function-Easily so, to find H(t) max you'd take this equation y= c- b^2/4a H(t) = 0- 29^2 / 4(-16) H(t) = -841/-64 H(t) = ? find out

OpenStudy (broskishelleh):

13.140625

OpenStudy (anonymous):

Yes, which rounds to 13.1

OpenStudy (broskishelleh):

Ye3